OpenStudy (anonymous):

How to solve an equation using factoring and quadratic formula ? x^+15x-54=0?

OpenStudy (anonymous):

X^2+18x-3x-54=0 x(x+18)-3(x+18)=0 x=3, -18

OpenStudy (anonymous):

For factorizing, Find two numbers whose product is -54 and sum is 15 18 & -3 \[x^2+18x-3x-54=0\] \[x(x+18)-3(x+18)=0\] \[(x+18)(x-3)=0\] x={-18,3}

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

For quadratic formula, ax^2+bx+c=0 \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] So it would be \[\frac{-15 \pm \sqrt{225+ 216}}{2}\] =6,-18

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

same equation, different roots..how come? its 3 and -18

OpenStudy (anonymous):

Uh sorry, typo, its 3,-18